Circular No. 6116

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET McNAUGHT-RUSSELL (1994u)

The following additional positions have been reported:

1994 UT α_{2000} δ_{2000} m_{1} Observer
Dec. 15.00328 7^{h}10^{m}54.^{s}29 −25^{o}08′19′′.0  Pravec
15.00635 7 10 54.10 −25 08 20.6  _{”}
15.01316 7 10 53.65 −25 08 24.2 16.3  _{”}
15.01656 7 10 53.48 −25 08 25.9  _{”}
15.59841 7 10 17.44 −25 14 12.7  Kojima
15.64446 7 10 14.43 −25 14 37.7  _{”}

P. Pravec (Ondřejov). 0.65-m f/3.6 reflector + R filter. m_{2} = 17.2; 0′.3 tail

in p.a. 340^{o}; comet only 15^{o} above horizon in very good sky conditions. T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6 reflector + CCD.

Communicated by S. Nakano, Sumoto.

Preliminary parabolic orbital elements by Nakano from 12 observations, Dec. 12–15:

T = 1994 Sept.10.868 TT  ω = 169.^{o}031
Ω = 221.569 }2000.0
q = 1.26647 AU  i = 30.948
1994/95 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 14 7^{h}11.^{m}93 −24^{o}57.^{′}6 1.125 1.874 125.^{o}3 25.^{o}4 16.5
19 7 06.70 −25 43.0 1.163 1.925 127.2 24.0 16.7
24 7 01.33 −26 10.9 1.206 1.976 128.8 22.8 16.9
29 6 56.02 −26 22.3 1.251 2.029 130.0 21.8 17.1
Jan. 3 6 50.95 −26 18.6 1.301 2.081 130.8 21.0 17.3
8 6 46.30 −26 01.3 1.354 2.134 131.2 20.3 17.5
13 6 42.19 −25 32.5 1.412 2.187 131.1 19.8 17.6
18 6 38.71 −24 53.9 1.473 2.241 130.5 19.5 17.8
23 6 35.91 −24 07.6 1.539 2.294 129.5 19.3 18.0

SUPERNOVA 1994ag IN ANONYMOUS GALAXY N. B. Suntzeff reports that his observations (details as given on IAUC

6115) on Dec. 12.2 UT show that the object reported on IAUC 6113 has features consistent with Na D and red TiO bands, suggesting that the object designated SN 1994ag is probably an early M dwarf.

1994 December 15 Daniel W. E. Green